OpenStudy (anonymous):

The equation of line AB is y = 5x + 1. Write an equation of a line parallel to line AB in slope-intercept form that contains point (4, 5).

OpenStudy (bibby):

What should the slope of the new line be?

OpenStudy (anonymous):

5x +1?

OpenStudy (bibby):

y=mx+b where m is the slope. I just want the m value of the new line

OpenStudy (anonymous):

5?

OpenStudy (bibby):

correct. now how do we find the equation of a line with slope 5 that passes through (4,5)?

OpenStudy (anonymous):

i have no idea ? is there a formula

OpenStudy (bibby):

there's something called point slope form. it looks like this \(y – y_1 = m(x – x_1)\) where x1 and y1 are the x and y coordinates of the point we want

OpenStudy (bibby):

god dammit y-y1=m(x-x1)

OpenStudy (anonymous):

haha ok so do i substitute y and x with (4, 5) ?

OpenStudy (bibby):

well x1 and y1*

OpenStudy (anonymous):

but how would i solve that i dont have a y or x?

OpenStudy (bibby):

you don't have to solve it. you just simplify it and then get it into slope-intercept form (y=mx+b)

OpenStudy (anonymous):

y- 5 = m (x -4)

OpenStudy (bibby):

correct. what is m?

OpenStudy (anonymous):

i don't know :( how do i get m?

OpenStudy (bibby):

m is what we call the slope. if we want parallel lines, we want the same slope as the original line

OpenStudy (anonymous):

5x + 1?

OpenStudy (bibby):

just 5. y=mx+b here the 5 is attached to the x, so m=5

OpenStudy (anonymous):

wait i dont understand my choices are A) y = 5x − 15 B) y = 5x + 15 C) y = 1 over 5x + 21 over 5 D)

OpenStudy (bibby):

y- 5 = 5 (x -4) distribute the 5 y-5 = 5x-20 now simplify

OpenStudy (anonymous):

y = 5x - 15

OpenStudy (bibby):

yep

OpenStudy (anonymous):

thankyou !!

OpenStudy (bibby):

np np

OpenStudy (anonymous):

Hi